OpenStudy (anonymous):

The areas of three faces of a right rectangular prism are 5, 7, and 11. What is its volume?

Directrix (directrix):

>Take the base to be the triangle of side 5, 7,11 If the solid is a right, rectangular prism, then its bases will be rectangles, not a triangles.

OpenStudy (anonymous):

Let x and y be the sides of the base and z the height. We have \[ x y = 5 x z = 7 y z = 11 ] We need V = xyz. Can you manage it from here?

OpenStudy (anonymous):

To find x and y, notice that by dividing the last two equations \[ \frac x y =\frac 7 {11}\\ x = \frac {7 y} {11} \] replace in the first equation, you get \[ x y = 5\\ \frac {7 y} {11} y= 5 \\ 7 y^2 = 55\\ y^2 = \frac {55}{11}\\ y= \frac { \sqrt {55} }{\sqrt{11}} \] Find x, then z . Your answer should be x y z

OpenStudy (anonymous):

\[ x = \frac 7 {11} y=\frac 7 {11} \frac { \sqrt {55} }{\sqrt{11}}=\frac { \sqrt {35} }{\sqrt{11}} \]

OpenStudy (anonymous):

\[ x = \frac 7 {11} \\y=\frac 7 {11} \frac { \sqrt {55} }{\sqrt{11}}\\ z= \frac {11}y= 11 \frac { \sqrt {11} }{\sqrt{55}}=\frac { \sqrt {77} }{\sqrt{5}}\\\\ \]

OpenStudy (anonymous):

Sqrt[35/11] Sqrt[55/7] Sqrt[77/5] Now compute \[ V = x y z=\sqrt { \frac {(35)(55)(77)} {(11)(7) (5)} }=\sqrt{385} \]

OpenStudy (anonymous):

A quickest way is to multiply the three original equations \[ x^2 y^2 z^2 = (5)(7)(11)= 385\\ x y z =\sqrt{385} \]